OpenStudy (anonymous):

a ceramics teacher divided 20 kg of clay as equally as possible among the 29 student in her class. about how much clay did each student get?

OpenStudy (ranga):

Divide 20 by 29.

OpenStudy (anonymous):

I try that it din not work

OpenStudy (ranga):

When you say it did not work what do you mean? Did they specify how many decimals they want the answer to have?

OpenStudy (anonymous):

there are 2.20462 pounds in a kg. 44.0924 pounds = 20 kg divide 44.0924 by 29 so about 1.52042758621 or 1.5 pounds per student

OpenStudy (ranga):

Did they say they want the answer in pounds?

OpenStudy (anonymous):

well the key word is about

OpenStudy (anonymous):

they want the answer in decimal can you help me with that please

OpenStudy (ranga):

20/29 = 0.6896551 If you round it off to: 1 decimal: 0.7 2 decimals: 0.69 3 decimals: 0.690 By about they mean just one decimal place the answer will be 0.7 kg per student

OpenStudy (anonymous):

so is the answer 0.7 ?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

thank you so much u a great help can I ask you some more question because people think am really dump

OpenStudy (anonymous):

i don't blame them i am dump

OpenStudy (ranga):

ask me one more.

OpenStudy (anonymous):

thanks

OpenStudy (ranga):

will u be typing the question here?

OpenStudy (anonymous):

juan worked 2.5 hours each day from monday through friday plus 7.5 hours on saturday. his total pay for the week was $128. how much did he earn per hour?

OpenStudy (anonymous):

so do you know the answer?

OpenStudy (anonymous):

ok can u explain more I kind of don't understand

OpenStudy (anonymous):

the answer is not $12.80

OpenStudy (ranga):

Are you doing the work online or from a textbook?

OpenStudy (anonymous):

my teacher print the pare out. but like in my answer choice there is not 12.80

OpenStudy (anonymous):

paper

OpenStudy (ranga):

unless you have not typed the number right in the problem the answer is 12.8. Can you type all the answer choices they have?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Juan worked 2.5 hours each day from Monday through Friday plus 7.5 hours on Saturday. his total pay for the week was $128. how much did he earn per hour?

OpenStudy (anonymous):

that the question

OpenStudy (ranga):

Oh, ok I got it now.

OpenStudy (ranga):

He worked 2.5 hours each day, Monday through Friday. Monday thru Friday is 5 days. So 5 x 2.5 = 12.5 hours Saturday he worked 7.5 hours. So total hours worked Monday through Saturday is: 12.5 + 7.5 = 20 hours. His total pay = $128 So his hourly wage = Total Pay / Total hours worked = 128 / 20 = 6.40 $6.40 / hour

OpenStudy (anonymous):

thank you that the answer

OpenStudy (ranga):

Just one more question.

OpenStudy (anonymous):

ok give me a second to type

OpenStudy (anonymous):

the 26-oz of chocolate fudge costs $3.12. the 16-oz jar costs $2.40. how much can you save on each ounce by choosing the better buy?

OpenStudy (ranga):

26 oz jar costs $3.12 Therefore, 1 oz costs 3.12/26 = $0.12 per ounce 16 oz jar costs $2.40 Therefore, 1 oz costs 2.40/16 = $0.15 per ounce So the better buy will be the one that costs less per ounce. So the difference is: $0.15 - $0.12 = $0.03 or 3 cents.

OpenStudy (anonymous):

thank you. you been a great help

OpenStudy (ranga):

you are welcome. good night.

OpenStudy (anonymous):

good night